# FAITH: A Framework for Assessing Intrinsic Tabular Hallucinations in Finance

### v1

@@ 1. Introduction
This paper focuses on intrinsic hallucinations—a form of LLM error where the generated output is inconsistent with the provided input context (Bang et al., 2025). Many financial tasks require precise extraction, summarization, calculation, interpretation, and reasoning based on structured inputs such as financial tables. Errors in these processes can significantly undermine decision-making. Intrinsic hallucinations are already concerning in such cases; when LLMs rely instead on web search results or internal knowledge rather than the actual financial input, the risk of hallucination becomes even greater. Undetected hallucinations can propagate through complex, automated systems, leading to misleading analyses, flawed investment strategies, or regulatory breaches.
@@
Yet, evaluating hallucinations in a systemic and robust approach is a non-trivial challenge, mainly due to the lack of finance-specific hallucination evaluation datasets. Most existing datasets are derived from general-domain texts such as Wikipedia (Gu et al., 2024; Bang et al., 2025), whereas financial applications often involve complex, numerically grounded data - including tables, charts, and document inputs - and require specialized reasoning and contextual understanding. More importantly, there is a clear need for an automated and scalable approach to creating evaluation datasets for hallucination detection in real-world financial settings. Manual annotation is resource-intensive and can hardly keep up with evolving LLM behaviors or the diversity of proprietary financial data. Thus, inspired by recent advancements (Yu et al., 2024; Luo et al., 2024; Hu et al., 2024), our study addresses this gap by proposing a rigorous and scalable methodology that uses actual financial annual reports—where ground truth values are explicitly known—to automatically construct benchmarking tasks for evaluating intrinsic hallucinations in financial LLM applications.
@@
In this study, we address a core challenge in financial LLM applications: accurately identifying, calculating, and presenting numerical values based on financial statements (input context). We introduce a novel context-aware masked span prediction task, where numerical values from real annual reports are masked and LLMs are prompted to recover them. To ensure reliable evaluation, we develop a novel filtering method that selects only those values with a unique, consistent, and answerable ground truth. A central contribution is our taxonomy of four mutually exclusive and exhaustive financial reasoning types: Direct Lookup, Comparative Calculation (The value is a function of the same variable in different periods, e.g., year-over-year growth), Bivariate Calculation (the value is a function of two variables, e.g., gross margin), and Multivariate Calculation (more complex dependencies). This classification enables structured evaluation across reasoning complexity. Using our approach, we build a large-scale benchmark dataset from S&P 500 annual reports and evaluate leading LLMs on their intrinsic hallucination rates. Results show that even top models frequently hallucinate, especially on complex tasks.
Our work provides an essential foundation for evaluating hallucinations in finance, a critical first step before effective mitigation. The main contributions are as follows:
@@ A.1. Financial Span Answerability Annotation
For each highlighted span substring (shown as <SPAN>{span}</SPAN>) in the sentence, determine if the span is: - unanswerable: Spans that - do not represent some type of numeric financial data (e.g., phone numbers, pincodes, or any other noise), or - the span cannot be derived from or supported by the table by any method. - answerable: Spans that - can be directly found in the table, or - can be derived through some calculations (such as addition, subtraction, multiplication, or division), or - can be inferred via deeper reasoning involving multiple table entries Instructions: (1) Carefully analyze every given table and the given sentence. (2) For each <SPAN>{span}</SPAN>, in the sentence: - Label the span as answerable or unanswerable (3) Provide a concise explanation for your reasoning. Output Format:
@@
"reasoning": <A detailed explanation for each highlighted
span in the sentence - if it is answerable: how it can
be matched, derived, or inferred from the table,
or else if it is unanswerable: a brief reason why.>
@@ A.2. Financial Metric Prediction
**Instructions:** (1) **Analyze the Request:** Carefully read the sentence and examine the provided tables to understand what information is needed to fill in the blank. (2) **Reason Step-by-Step:** Before providing the final answer, write out your reasoning process. Explain how you will find the value, including any calculations. - If you are extracting a value, mention which table and cell you are getting it from. - If you are performing a calculation, show the formula and the values you are using. (3) **Categorize Your Reasoning:** Based on your reasoning, classify it into one of the following scenarios: - **A. Direct Lookup:** The value is directly extracted from a single cell in a table. - **B. Simple Calculation (Single Metric):** The result is calculated from a single metric across different time periods, categories, or rows (e.g., calculating a year-over-year change). - **C. Simple Calculation (Two Metrics):** The result is calculated using two different metrics (e.g., calculating a ratio). - **D. Complex Calculation:** The reasoning involves more than two metrics or multiple complex steps. (4) **Format the Final Answer:** After your reasoning, provide the final answer in a JSON block with the following structure. **JSON Output Format:**
@@
"necessary_metrics": ["<metric_name_1>",
"<metric_name_2>", ...],
"reference": ["<table_identifier_1>",
"<table_identifier_2>", ...]
**Field Explanations:**
‘answer‘: The value to fill in the masked blank. Format it professionally (i.e., with rounding and units, etc.).
‘scenario‘: One of ”A”, ”B”, ”C”, or ”D” based on your reasoning.
‘necessary_metrics‘: A list of all metric names from the tables required to derive the answer.
‘reference‘: A list of all table identifiers for the tables used.
**Inputs:**
**Tables:** {tables_with_pretext}
**Sentence:** {pre_sentence} {sentence} {post_sentence}

### v2

@@ 1. Introduction
Recent empirical studies have begun to reveal that hallucinations in finance exhibit domain-specific characteristics compared with those observed in general-domain NLP benchmarks. For instance, Shah et al. (2025) find that LLMs’ factual accuracy on historical data varies across company size and financial periods, while Kang and Liu (2023) and Sarmah et al. (2024) demonstrate that even state-of-the-art models frequently misinterpret financial terms or misreport numerical values. Complementary work such as Roychowdhury (2024) outlines stages toward hallucination-minimized systems for financial decision-makers. Together, these studies underscore the urgency of domain-specific evaluation frameworks that capture the unique reasoning and numerical dependencies underlying financial data. However, a systematic, fine-grained benchmark that directly measures hallucinations in finance remains lacking.
@@
This paper focuses on intrinsic hallucinations, a form of LLM error where the generated output contradicts or misrepresents the input financial context (Bang et al., 2025). Many financial tasks require precise extraction, summarization, calculation, and reasoning based on structured statements such as balance sheets or income tables. Errors in these processes can significantly undermine analytical integrity. While intrinsic hallucinations are already concerning in such settings, reliance on external knowledge or web retrieval amplifies risks. Undetected hallucinations can propagate through automated reporting pipelines or investment algorithms, potentially resulting in misleading insights, compliance violations, or financial losses. From a practical standpoint, even modest hallucination rates—such as 10–20% in complex financial reasoning—can have outsized impacts when scaled across high-stakes decisions in portfolio management or regulatory reporting.
@@
Yet, evaluating hallucinations in a systematic and robust manner remains a non-trivial challenge due to the scarcity of finance-specific benchmarks. Existing datasets are predominantly derived from general-domain texts such as Wikipedia (Gu et al., 2024; Bang et al., 2025), whereas financial applications involve numerically grounded, context-dependent data and specialized terminology. More importantly, there is a clear need for an automated and scalable approach to creating evaluation datasets for hallucination detection in real-world financial settings. Manual annotation is resource-intensive and can hardly keep up with evolving LLM behaviors or the diversity of proprietary financial data.
@@
To address these gaps, we develop FAITH (A Framework for Assessing Intrinsic Tabular Hallucinations in Finance) - an automated and scalable approach for constructing intrinsic hallucination evaluation tasks directly from real financial annual reports, where ground-truth numerical values are explicitly available. We address a core challenge in financial LLM applications: accurately identifying, calculating, and presenting numerical values based on financial statements (input context). We introduce a novel context-aware masked span prediction task, where numerical values from real annual reports are masked and LLMs are prompted to recover them. To ensure reliable evaluation, we develop a novel filtering method that selects only those values with a unique, consistent, and answerable ground truth. A central contribution is our taxonomy of four mutually exclusive and exhaustive financial reasoning types: Direct Lookup, Comparative Calculation (The value is a function of the same variable in different periods, e.g., year-over-year growth), Bivariate Calculation (the value is a function of two variables, e.g., gross margin), and Multivariate Calculation (more complex dependencies). This classification enables structured evaluation across reasoning complexity. Using our approach, we build a large-scale benchmark dataset from S&P 500 annual reports and evaluate leading LLMs on their intrinsic hallucination rates.
Empirically, our results reveal a stratified hierarchy of model reliability in financial reasoning. Proprietary frontier models such as Claude-Sonnet-4 and Gemini-2.5-Pro achieve high overall accuracy, yet even these systems exhibit 10–20% error rates on multi-step numerical reasoning. In contrast, smaller open-source models often fail catastrophically on complex calculations, scoring near zero in multivariate scenarios. This degradation with reasoning complexity highlights an urgent limitation of current architectures—the difficulty of maintaining factual consistency across chained numerical operations. Importantly, such residual hallucination rates, though seemingly modest, could translate into substantial financial misjudgments in investment analytics, credit risk assessment, or regulatory compliance when scaled to production systems.
Our main contributions are as follows:
@@ 6. Conclusion and Future Work
7. Limitation
FAITH provides a systematic framework for assessing intrinsic tabular hallucinations in financial contexts, yet several limitations remain.
Our benchmark draws on public U.S. company filings and curated financial tables. This design ensures data reliability and comparability but limits coverage of diverse reporting styles, regional accounting standards, and sector-specific terminologies. Nevertheless, the framework is readily expandable to other companies, industries, and jurisdictions, provided that structured financial data are available. Future work can extend FAITH to international filings and more unstructured financial disclosures.
When constructing masked instances, we retain only those that three frontier LLMs unanimously identify as answerable. Although a pilot study validates the reliability of this heuristic, it may systematically exclude the most challenging or ambiguous spans—precisely those that elicit disagreement or uncertainty across models. As a result, FAITH may slightly underestimate the intrinsic difficulty of hallucination detection. A systematic examination of these ambiguous, inter-model disagreement cases constitutes a valuable direction for future research, potentially enhancing both dataset representativeness and evaluation robustness.
@@ A.1. Financial Span Answerability Annotation
You are given the following tables from a 10-K annual report:
{tables_with_pretext}
@@
For each highlighted span substring (shown as [span]) in the sentence, determine if the span is:
unanswerable: Spans that
do not represent some type of numeric financial data (e.g., phone numbers, pincodes, or any other noise), or
the span cannot be derived from or supported by the table by any method.
answerable: Spans that
can be directly found in the table, or
can be derived through some calculations (such as addition, subtraction, multiplication, or division), or
can be inferred via deeper reasoning involving multiple table entries
Instructions:
Carefully analyze every given table and the given sentence.
For each span in the sentence, label the span as answerable or unanswerable.
Provide a concise explanation for your reasoning.
Output Format:
@@
"reasoning": "<A detailed explanation for each highlighted
span in the sentence if it is answerable: how it can
be matched, derived, or inferred from the table; or if
it is unanswerable: a brief reason why.>",
@@ A.2. Financial Metric Prediction
Instructions:
Analyze the Request: Carefully read the sentence and examine the provided tables to understand what information is needed to fill in the blank.
Reason Step-by-Step: Before providing the final answer, write out your reasoning process. Explain how you will find the value, including any calculations.
If you are extracting a value, mention which table and cell you are getting it from.
If you are performing a calculation, show the formula and the values you are using.
Categorize Your Reasoning: Based on your reasoning, classify it into one of the following scenarios:
A. Direct Lookup: The value is directly extracted from a single cell in a table.
B. Simple Calculation (Single Metric): The result is calculated from a single metric across different time periods, categories, or rows (e.g., calculating a year-over-year change).
C. Simple Calculation (Two Metrics): The result is calculated using two different metrics (e.g., calculating a ratio).
D. Complex Calculation: The reasoning involves more than two metrics or multiple complex steps.
Format the Final Answer: After your reasoning, provide the final answer in a JSON block with the following structure.
JSON Output Format:
@@
"necessary_metrics": ["<metric_name_1>", "..."],
"reference": ["<table_identifier_1>", "..."]
Field Explanations:
answer: The value to fill in the masked blank. Format it professionally (i.e., with rounding and units, etc.).
scenario: One of A”, B”, C”, or D” based on your reasoning.
necessary_metrics: A list of all metric names from the tables required to derive the answer.
reference: A list of all table identifiers for the tables used.
Inputs:
Tables: {tables_with_pretext}
Sentence: {pre_sentence} {sentence} {post_sentence}